Early in its fiscal year ending December 31, 2018, San Antonio Outfitters finalized plans to expand operations. The first sta1. Determine the initial values of the various assets that San Antonio acquired or constructed during 2018. The company usesnterest expense $ 112,830 *
Answer #1
Please hit LIKE button if this helped. For any further explanation, please put your query in comment, will get back to you.
Requirement 1
Land
Purchase price (determined below) $                                                                 841,986
Closing costs $                                                                   28,000
Removal of old building $                                                                   78,000
Clearing and grading $                                                                   58,000
$                                                              1,005,986
Purchase price of land:
      Cash paid $                                                                 280,000
      Value of note (From below working) $                                                                 561,986
$                                                                 841,986
† Present value of note payment:
PV = $680,000 (.82645) $                                                                 561,986
Present value of $1: n = 2, i = 8% (from Table 2)
Land improvements
Parking lot and landscaping $                                                                 325,000
Building
Construction expenditures:  
5-1 $                                                              2,400,000
7-1 $                                                              1,900,000
9-1 $                                                              1,380,000
10-1 $                                                              2,280,000
    Total expenditures $                                                              7,960,000
Interest capitalized (determined below) $                                                                 209,500
      Total cost of building $                                                              8,169,500
Average accumulated expenditures:
May 31, 2018 $2,400,000 x   6/6 = $                 2,400,000
July 30, 2018 1,900,000 x   3/6 = $                    950,000
September 1, 2018 1,380,000 x   2/6 = $                    460,000
October 1, 2018 2,280,000 x   1/6 = $                    380,000
$                 4,190,000
Interest capitalized:
$4,190,000 x10% x 6/12 = $                                                                 209,500
Equipment and furniture and fixtures
      Initial
Percent of Total     Valuation
Fair Value Fair Value % x $680,000
Equipment $                                                                 585,000                               0.75 $           510,000
Furniture & fixtures $                                                                 195,000                               0.25 $           170,000
Totals $                                                                 780,000                               1.00 $           680,000
Initial valuation:
Equipment $                                                                 510,000
Furniture & fixtures $                                                                 170,000
Requirement 2
Interest expense:
Note issued to purchase land and building,
     $561,986 x 10% x 9/12 = $                                                                   42,149
Construction loan, $4,200,000 x 10% x 8/12 $                                                                 280,000
Long-term note, $2,800,000 x 10% $                                                                 280,000
Long-term bonds, $4,800,000 x 7% $                                                                 336,000
      Total $                                                                 938,149
Less: Interest capitalized (determined above) $                                                                -209,500
Interest expense $                                                                 728,649